OpenStudy (anonymous):

Write an equation that passes through the given point and that is perpendicular to the given line in slope intercept form (-2,-4), y=-2/7x+1

OpenStudy (mathstudent55):

Perpendicular lines have slopes that are negative reciprocals. That means that when you multiply together the slopes of perpendicular lines you get -1. In a practical way, if you have the slope of a line and want the slope of a perpendicular, write the slope as a fraction, then flip the fraction, then change the sign.

OpenStudy (mathstudent55):

First let's find the slope of the given line: y = -(2/7)x + 1 Since this equation is already written in y = mx + b form, we can read the slope directly. m = -2/7

OpenStudy (anonymous):

ok thanks so then what if the slope was already was negative,will that mean that the slope will turn into positive?

OpenStudy (mathstudent55):

Correct. Now we figure out the slope of a perpendicular to this line: m1 = -2/7; m2 = 7/2

OpenStudy (mathstudent55):

We flip the fraction and change sign. The slope of the perpendicular is 7/2. Now our question is what is the equation of the line that has slope 7/2 and passes through (-2, -4)?

OpenStudy (mathstudent55):

You can use the slope-intercept equation, y = mx + b, let m = 7/2, x = -2, and y = -4, and solve for b

OpenStudy (anonymous):

Yes so then when we plug in the X and Y of the given points it will be 4=7/2(-2)+b

OpenStudy (mathstudent55):

Negative 4, not positive 4

OpenStudy (anonymous):

Oh yes my bad.the slope will be b=3 ?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

So then the new intercept form will be y=7/2x+3

OpenStudy (mathstudent55):

Correct

OpenStudy (anonymous):

So then the only thing different than the parallel line equations is changing the slope into a reciprocal form right?

OpenStudy (mathstudent55):

Right. Figuring out the new slope is different, but then getting the equation is the same procedure.

OpenStudy (anonymous):

Can you help me in another problem in another post please?